7. A countrys economy in 2015 and 2016 is summarized in the following table. You must show all your work 015 2016 Price Turmeric (domestically produced and consumed) Cardamom (exported) Sesame seeds Price 1000$20 $10 100 $15 2000 100 1200 2200 20 a) Assuming that the countrys population did not change between 2015 and 2016, what happened to real GDP per capita between the two years? Explain your answer (1 point) b) Calculate the inflation rate experienced by the economy between the two years, using GDP deflators and 2015 as base year. (1 point) 8. The table below shows the annual incomes of the only residents of two countries A and B. Draw each countrys Lorenz curve and determine which country has a more unequal income distribution. Person Incomes in C Incomes in C 20 25 1000 3050 40 50 2000 60 100 7&8 please
Answer #1
item price (2015 ) quantity (2015 ) value(2015 ) price (2016 ) quantity (2016 ) value (2016 )( at 2015 prices )
turmeric 15 1000 15000 20 1200 18000
cardamom 5 2000 10,000 10 2200 11000
sesame seeds 50 100 5000 100 20 1000

real gdp 2015 = turmeric value + cardamom value - sesame seeds value

= 15000 + 10,000 - 5000

= $20,000

real gdp 2016 = 18000 + 11000 - 1000

= $28000

since population ha remained same , the Real gdp per capita has increased .

b ) gdp deflator (2015 ) nominal gdp *100 / real gdp

= 20,000 *100/ 20,000 = 100

gdp deflator (2016 ) = nominal gdp 2016 *100 / real gdp

= 44000*100/28000 = 157.14

inflation rate = gdp deflator 2016 - gdp deflator 2015 /100

= 157.14 - 100/100

.57 or 57.14%

( done 1 question as per policy )